OpenStudy (anonymous):

If one cup is equal to 8 fluid ounces, how many fluid ounces are in 3 cups? Numerical Answers Expected! Answer for Blank 1: WILL FAN N GIVE U A MEDAL!!!<<<<

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

1 cup = 8 oz 3 * 1 cup = 3* 8 oz = 24 oz